Construction company X signed a contract to build a hotel for Donald at the price of HKD 720 million. The cost of the construction at the time of signing the contract was expected to be HKD 900 million with 20% probability, HKD 800 million with 30% probability, HKD 600 million with 30% probability, and HKD 500 million with 20% probability. Donald values the hotel at HKD 850 million.

  1. How much should X pay to Donald in case of a breach under the expectation damage? (5 points)
  2. What is the likelihood of the breach under the expectation damage?  Is this frequency of a breach efficient? (5 points)
  3. Suppose that the court cannot assess the value to the Donald well and the court only can assess the Donald’s valuation as HKD 750 million (and both X and Donald knows about this) in calculating the expectation damage.  Also suppose that the cost of renegotiating contract is very high. What is the likelihood of a breach in this case?  Is this frequency efficient? (8 points)
  4. Consider the same situation as above except that the cost of (re)negotiation is zero. a) If the realized cost were HKD 900 million, what would you expect to happen? b) What if the realized cost were HKD 800 million? (8 pints, each)
  5. Discuss your answer in the above questions (I mean, questions 3 and 4) by relating it to the Coase theorem (8 points)
Answer #1

The Coase theorem states that " In the absence of transaction costs, if property rights are well defined & tradable, then voluntary negotiations will lead to efficiency. The Coase theorem works on the principle that though the rights can have a different association with regards to the initial pay off , the same can be renegotiated until they are allocated efficiently or both the parties to the dispute have reached to an agreeable conclusion.

From the problem which is stated above, we find that a contract is being signed by Construction Company X with Donald and there are various costs associated with it both initially and the probable risk percentage associated with it. We have to analyse the situation and find the expectation damage and frequency of breach.

An expectation damage occurs when either of the parties to the contract has a cost over-run or when one has to spend more than the expected costs. Let us calculate the expected cost and observed cost and estimate the damage as per the problem with all the cost ( in millions )

Initial cost( in millions) at the time of signing the contract = 720

Expected Cost with 20% probability = 900 or 500

Expected Cost with 30% probability = 800 or 600

If the court negotiates with a value of 750 , then expected damage on the part of Donald = 850-750= 100 million, in a similar way, the expected damage on the part of construction company X = 750-720= 30 million. The frequency of breach is not efficient as Donald suffers an expected loss 3 times more than Construction Company X.

If the realized cost is 900 million which has 20% probability, then actual cost = 900- (900x 0.2) = 900-180 = 720 million, which was the initial price of contract. The expected damage in this case to Donald is nill .

If the realized cost is 800 million which has 30% probability, then actual cost = 800 - (800 x0.3) = 800-240= 560 million, which is less than the initial price of the contract. The expected damage to Donald in this case is more, but less to the construction company as Donald is paying 720 million for the contract.